Alice leaves her house, driving east at 45 miles per hour (mph). Thirty minutes later, her husband Dave notices she forgot her cell phone and sets off after her. How fast must Dave travel in order to catch up with Alice 3 hours after he leaves
Answer by stanbon(75887)

Alice leaves her house, driving east at 45 miles per hour (mph). Thirty minutes later, her husband Dave notices she forgot her cell phone and sets off after her. How fast must Dave travel in order to catch up with Alice 3 hours after he leaves
----------------------
Alice DATA:
time = (7/2) hr ; rate = 45 mph ; distance = rt = (7/2)45 miles
--------------------------
Dave DATA:
time = 3 hr ; distance = (7/2)45 = 157.5 miles ; rate = x mph
---
Equation:
Dave's rate = dist/time = 157.5/3 = 52.5 mph
=============================================
Cheers,
Stan H.
=============================================